OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

Equation form: y= mx+ b You got m= -6: y= -6x+ b y_int of 9-----> passes the point (0,9) now you have: 9=-6(0)+b 9= 0+b b=9. Therefore the equation is: y= -6x +9

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (anonymous):

You are welcm!